OpenStudy (anonymous):

will medal Use the distributive property to rewrite the expression and solve. 6 94 = 6___+ 6 4 =___

OpenStudy (gabebae):

wait what? looks confusing

OpenStudy (anonymous):

oh hold on

OpenStudy (anonymous):

the equation is 6*94=6*___+6*4=___

OpenStudy (anonymous):

dots didn't show up, sorry

OpenStudy (anonymous):

better?

OpenStudy (gabebae):

hold on I was afk

OpenStudy (anonymous):

u back @GabeBae?

OpenStudy (gabebae):

yup

OpenStudy (anonymous):

Use the distributive property to rewrite the expression and solve. 6*94=6*___+6*4=___

OpenStudy (gabebae):

I got 6*50+ 4*66 = 564 94*6 is 564

OpenStudy (anonymous):

what do i put in the blanks?

OpenStudy (gabebae):

50 and 66

OpenStudy (anonymous):

it said incorrect

OpenStudy (gabebae):

I don't know what it was trying to ask